OpenStudy (anonymous):

Could anyone please help me with this calculus problem... Find the limit: limx--->0 (1-sin^2x/x)

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}(\frac{ 1-\sin ^{2}x }{ x } )\]

OpenStudy (anonymous):

So far I think that I would have to change the 10-sin^2x into cos^2... right ?

OpenStudy (anonymous):

*1-sin^2x*

OpenStudy (anonymous):

It won't change the fact that you have a nonzero number in the numerator and zero in the denominator. That is undefined or "does not exist".

OpenStudy (anonymous):

If you ever get a zero in the numerator and the denominator, then you may be able to manipulate the function into something with an answer. But 1/0 is simply a d.n.e.

OpenStudy (anonymous):

What do you mean by a non zero number ? are you referring to the 1 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

If you get 0/0, it is considered indeterminate and you might find an answer another way. But 1/0, an you are done.

OpenStudy (anonymous):

So you're saying that when this problem is solved it is a 1/0?

OpenStudy (anonymous):

Yes, put zero in for x.\[\frac{1-\sin ^{2}(0)}{0}=\frac{1}{0}\]

OpenStudy (anonymous):

Always start by just plugging the number in ("substitution"). If you get 0/0, try to manipulate then plug in again.

OpenStudy (anonymous):

okay, thank you for the help (:

OpenStudy (anonymous):

You're welcome.